OpenStudy (anonymous):

Hey all! Can someone help me out with my studyplan? I want to get this! How do I write 64^-2/3= in radical notation? And then evaluate the expression?

OpenStudy (anonymous):

\[a ^{\frac{ m }{ n }} = \sqrt[n]{a ^{m}}\] that's a power law. Try that.

OpenStudy (anonymous):

:( thanks but, not it.

OpenStudy (jdoe0001):

\(\bf a^{\frac{n}{m}} = \sqrt[m]{a^n}\\ a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (anonymous):

What do you mean can't you do...\[64^{\frac{ -2 }{ 3 }} = \sqrt[3]{64^{-2}}\]

OpenStudy (jdoe0001):

\(\huge a^{-\frac{n}{m}} =\cfrac{1}{\sqrt[m]{a^n}} \)

OpenStudy (anonymous):

oh wait - I think I put it in wrong...sorry..I hate this

OpenStudy (anonymous):

hmmm still not taking it

OpenStudy (anonymous):

Alright try this \[\frac{ 1 }{ \sqrt[3]{64^{2}} }\]

OpenStudy (anonymous):

ugh, it says to write it in radical notation - do not evaluate

OpenStudy (anonymous):

simplified I get a/64

OpenStudy (anonymous):

I did write it In radical form... I didn't simplify. Do you know what radical form is?

OpenStudy (anonymous):

the sq rt symbol

OpenStudy (anonymous):

I do not know why it will not tajke it

OpenStudy (anonymous):

Be careful and try again.

OpenStudy (anonymous):

I got it..I wasn't putting the power of 2 in....thank you so much!

OpenStudy (anonymous):

good job well done.!

OpenStudy (anonymous):

Thanks again - this study plan is killing me.